# Jackknife Empirical Likelihood Method for U Statistics Based on Multivariate Samples and its Applications

### v1

@@
Abstract
Empirical likelihood (EL) and its extension via the jackknife empirical likelihood (JEL) method provide robust alternatives to parametric approaches, in the contexts with uncertain data distributions. This paper explores the theoretical foundations and practical applications of JEL in the context of multivariate sample-based U-statistics. In this study we develop the JEL method for multivariate U-statistics with three (or more) samples. This study enhance the JEL methods capability to handle complex data structures while preserving the computation efficiency of the empirical likelihood method. To demonstrate the applications of the JEL method, we compute confidence intervals for differences in VUS measurements which have potential applications in classification problems. Monte Carlo simulation studies are conducted to evaluate the efficiency of the JEL, Normal approximation and Kernel based confidence intervals. These studies validate the superior performance of the JEL approach in terms of coverage probability and computational efficiency compared to other two methods. Additionally, a real data application illustrates the practical utility of the approach. The JEL method developed here has potential applications in dealing with complex data structures.
Empirical Likelihood,
Jackknife,
U Statistics,
Kernel,
Volume under ROC surface,
keywords:
[…]
1 Introduction
Empirical likelihood (EL) has emerged as a powerful tool in statistical inference, particularly valuable when traditional parametric assumptions about data distribution are either unknown or do not hold. EL operates on the principle of maximizing a nonparametric likelihood function constrained by observed data to perform hypothesis tests and construct confidence intervals. Owen’s papers ([15], [16]) have been instrumental in the development and popularization of EL. Furthermore, [17] highlights EL’s ability to automatically shape confidence regions, integrate additional constraints, and preserve the range and transformations inherent in the data. This flexibility and robustness have positioned EL as a cornerstone in modern statistical methodology, offering advantages in scenarios ranging from incomplete data to complex multivariate analyses ([19]).
The jackknife empirical likelihood (JEL) method is an extension of Owen’s foundational work that combines the EL method with the jackknife resampling technique. [7] introduced the JEL method as a modification of EL to deal with non-linear statistics. They illustrated this method using one- and two-sample U-statistics, paving the way for its application in diverse settings. [26] considered testing the equality of two multivariate means using the JEL method. [11] proposed a JEL-based non-parametric inferential procedure for non-smooth estimating functions and demonstrated its application. [18] extended the EL method to general estimating equations based on U-statistics and proposed JEL for vector U-statistics. Using this approach, they developed EL ratio tests for moment-based distribution characteristics. [20] developed a consistent JEL-based test for diagonal symmetry based on energy statistics.
Motivated by these studies, we generalize the JEL method to accommodate $q$-dimensional three (or more)-samples U-statistics. This generalization is needed for a versatile statistical tool capable of handling complex data structures and multiple sample comparisons without relying on distributional assumptions. This extension broadens the applicability of EL and JEL beyond traditional univariate and two-sample scenarios. Additionally, we explain the adaptation of JEL to $q$-dimensional $k$-sample based U-statistics, contribute to a deeper understanding of its theoretical foundations and practical implications in multivariate statistical analysis.
To give an application of described JEL method for three sample based U statistics, we applied it to a practical problem by constructing confidence intervals for the difference in volumes under surfaces (VUS) for two dependent variables. The VUS, introduced by [14], expands upon area under the receiver operating characteristic (ROC) by accommodating three-class diagnostic scenarios (a detailed discussion about VUS is provided in Section 2.1). For an application of the JEL method for $k$-sample-based U statistics, we construct confidence intervals for the difference of two dependent hypervolumes under the ROC manifolds (HUM). ROC curves are generalized to ROC manifolds for multi-category classification, and the analogous extension of the area under the ROC manifold is termed the HUM. For more details on HUM, refer to [10] and Chapter 3 of [25] (a detailed discussion about HUM is provided in Section 5.1).
The paper is organized as follows: In Section 2, we generalize JEL to $q$-dimensional, three-sample-based U-statistics. To provide an application of the described method, we consider the problem of constructing confidence intervals for difference of two dependent VUSs. We use the methods: JEL, U-statistics with asymptotic normality, and the kernel bootstrap to obtain the confidence intervals. Section 3 conducts extensive simulation studies to evaluate effectiveness of the confidence intervals for the difference of the VUSs under the Marshall-Olkin bivariate exponential and Bivariate Pareto models. In Section 4, we consider a real data analysis to demonstrate the findings. In Section 5, we further generalize the JEL U-statistics-based approach to $k\,(>3)$ multivariate sample U-statistics and illustrated the method using HUM.
2 JEL for three multivariate sample
@@
Let $\left(\mathbb{X}_{1},\ldots,\mathbb{X}_{n_{1}}\right),\left(\mathbb{Y}_{1},\ldots,\mathbb{Y}_{n_{2}}\right)$ and $\left(\mathbb{Z}_{1},\ldots,\mathbb{Z}_{n_{3}}\right)$ represent three independent random samples from populations with independent $q$-dimensional distribution functions $F$, $G$ and $H$, respectively, where $\mathbb{X}_{i}=(X_{11},X_{12},\dots,X_{1q})^{T},\;i=1,2,...,n_{1}$, $\mathbb{Y}_{j}=(Y_{11},Y_{12},\dots,Y_{1q})^{T},\;$ $j=1,2,...,n_{2},$ and $\mathbb{Z}_{k}=(Z_{11},Z_{12},\dots,Z_{1q})^{T},\;k=1,2,...,n_{3}$, and $q\geq 2$.
@@
Let $\theta=E[h\left(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}}\right)]\in\Re$ be the parameter of interest. A three-sample U-statistics of degree $\left(m_{1},m_{2},m_{3}\right)$ with a symmetric kernel $h$ is given by
@@
$U_{n_{1},n_{2},n_{3}}=$ $\frac{1}{{n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}\choose m_{3}}}\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ 1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{2}\\ 1\leq k_{1}<\ldots<k_{m_{3}}\leq n_{3}\end{subarray}}h\left(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}}\right),$ (1)
@@
By using the central limit theorem (see [6]), we have
[…]
Consider
(1) the U statistics based on all observations $U_{n_{1},n_{2},n_{3}}$, as defined by (1);
$U_{n_{1}-1,n_{2},n_{3}}^{(-i)}=\frac{1}{{n_{1}-1\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}\choose m_{3}}}\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}-1\\ i_{l}\neq i,\;l=1,2,\dots,m_{1},\\ 1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{2}\\ 1\leq k_{1}<\ldots<k_{m_{3}}\leq n_{3}\end{subarray}}\hskip-14.45377pth\left(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}}\right);$ (2)
$U_{n_{1},n_{2}-1,n_{3}}^{(-j)}=\frac{1}{{n_{1}\choose m_{1}}{n_{2}-1\choose m_{2}}{n_{3}\choose m_{3}}}\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ 1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{2}-1\\ j_{l}\neq j,\;l=1,2,\dots,m_{2},\\ 1\leq k_{1}<\ldots<k_{m_{3}}\leq n_{3}\end{subarray}}\hskip-14.45377pth\left(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}}\right);$ (3)
$U_{n_{1},n_{2},n_{3}-1}^{(-k)}=\frac{1}{{n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}-1\choose m_{3}}}\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ 1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{2}\\ 1\leq k_{1}<\ldots<k_{m_{3}}\leq n_{3}-1\\ k_{l}\neq k,\;l=1,2,\dots,m_{3}\end{subarray}}\hskip-14.45377pth\left(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}}\right).$ (4)
The jackknife pseudo-values for individual samples are defined as,
$V_{i,0,0}=n_{1}U_{n_{1},n_{2},n_{3}}-\left(n_{1}-1\right)U_{n_{1}-1,n_{2},n_{3}}^{(-i)},\;\;\forall\;\;i=1,2,\dots,n_{1},$ (5)
$V_{0,j,0}=n_{2}U_{n_{1},n_{2},n_{3}}-\left(n_{2}-1\right)U_{n_{1},n_{2}-1,n_{3}}^{(-j)},\;\;\forall\;\;j=1,2,\dots,n_{2}$ (6)
and
$V_{0,0,k}=n_{3}U_{n_{1},n_{2},n_{3}}-\left(n_{3}-1\right)U_{n_{1},n_{2},n_{3}-1}^{(-k)},\;\;\forall\;\;k=1,2,\dots,n_{3}.$ (7)
The jackknife pseudo-values based estimator of $\sigma^{2}$, proposed by [2] (also see [25], [4] and [1]), given as
$+\frac{1}{n_{3}\left(n_{3}-1\right)}\sum_{k=1}^{n_{3}}\left(V_{0,0,k}-\bar{V}_{0,0,\cdot}\right)^{2},$ (8)
where $V_{i,0,0}$, $V_{0,j,0}$ and $V_{0,0,k}$ are defined in equations (5), (6) and (7), respectively.
Now, we provide the asymptotic properties of the U-statistics $U_{n_{1},n_{2},n_{3}}$ using the estimator $\hat{\sigma}^{2}$. The following notations are required to describe the theorem as stated below.
@@
(b) Suppose that $\sigma_{100}^{2},\sigma_{010}^{2},\sigma_{001}^{2}>0$ holds. Let $S_{n_{1},n_{2},n_{3}}^{2}=\frac{m_{1}^{2}\sigma_{100}^{2}}{n_{1}}+\frac{m_{2}^{2}\sigma_{010}^{2}}{n_{2}}+\frac{m_{3}^{2}\sigma_{001}^{2}}{n_{3}},$ then,
[…]
and $\;\;\hat{\sigma}^{2}-S_{n_{1},n_{2},n_{3}}^{2}=o_{p}\left(\min\{n_{1},n_{2},n_{3}\}^{-1}\right).$
A two sided $(1-\alpha)$ level CI for $\theta$ based on the above theorem is
$\left(U_{n_{1},n_{2},n_{3}}-z_{\alpha/2}\hat{\sigma},U_{n_{1},n_{2},n_{3}}+z_{\alpha/2}\hat{\sigma}\right),$
@@
For proofs of parts (a) and (b) of the above theorem refer to pages 151-153 of [9]. The proof of the second part of (b) can be easily obtained using Chebyshev’s inequality and is thus omitted (see [4]).
The normal approximation method described above is computationally expensive. Additionally, [7] demonstrated that the JEL method, based on one and two sample U-statistics, outperforms the normal approximation method. Therefore, we develop the JEL method for the above problem to make inferences about the parameter $\theta$. To implement the JEL method, we combine all samples and treat them as a single sample. For this purpose, let $n=n_{1}+n_{2}+n_{3}$ and define $\mathbb{W}_{l}$ as

### v2

# Jackknife Empirical Likelihood for Multivariate three-Sample U-Statistics and its Applications

@@
Abstract.
We develop a jackknife empirical likelihood (JEL) framework for inference on parameters defined through multivariate three-sample U-statistic. From three independent multivariate samples, we construct JEL ratio statistic based on suitable jackknife pseudo-values and, under mild regularity conditions, establish a Wilks-type result showing that the log JEL ratio converges in distribution to a chi-square limit. This provides asymptotically valid confidence intervals for the parameter of interest without explicit variance estimation or heavy resampling. To illustrate the usefulness of the proposed method, we construct confidence intervals for differences in volume under the surface (VUS) measures, which are widely used in classification problems. Through Monte Carlo simulations, we compare the performance of JEL-based confidence intervals with those obtained from normal approximation of U-statistic and kernel-based methods. The findings indicate that the proposed JEL approach outperforms existing methods in terms of coverage probability and computational efficiency. Finally, we apply our methods to a recent real dataset.
Keyword: Jackknife empirical likelihood (JEL); Volume under surface (VUS); U-statistics.
$\dagger$}Corresponding Author mail: skkattu@isichennai.res.in
1. Introduction
Empirical likelihood (EL) has emerged as a useful tool in statistical inference when traditional parametric assumptions about the distribution of the data are either unknown or do not hold. EL operates on the principle of maximizing a nonparametric likelihood function with constraints on observed data with applications to hypotheses testing and confidence interval estimation. Early work on EL is due to Owen (owen1988empirical, owen1990empirical). owen2001empirical highlighted that empirical likelihood (EL) inherently constructs confidence regions, accommodates supplementary constraints, and preserves the natural scale and transformations of the data. This adaptability has established EL as a useful methodology in in handling incomplete data and addressing complex multivariate problems (jian2008).
For nonlinear statistics such as U-statistics, however, the original EL formulation can be difficult to implement directly, because the constraints involve nonlinear functionals of the empirical distribution. To address this difficulty, jing2009jackknife proposed the jackknife empirical likelihood (JEL) method, which constructs empirical likelihood ratios based on jackknife pseudo-values. Their work established Wilks-type results for one- and two-sample univariate U-statistics and demonstrated that JEL often improves coverage accuracy over normal approximation methods while avoiding explicit variance estimation. Since then, JEL has been extended to various settings, including multivariate mean comparisons Wang2013, non-smooth estimating equations Li2016, and tests based on energy statistics SANG2020, underscoring its flexibility for complex nonparametric functionals. sang2021jackknife developed a jackknife empirical likelihood procedure for nonparametric $k$-sample tests based on categorical Gini correlations, and established a chi-square limiting distribution for the resulting test statistic.
Despite this progress, existing JEL developments for U-statistics are predominantly limited to one- or two-sample problems, or to vector parameters in a single-sample setting. In many applications, however, the parameter of interest is naturally expressed as a functional of multivariate three-sample U-statistics. Examples include ordering probabilities in three-arm clinical trials, multivariate stress–strength reliability problems, and measures of three-class separation in diagnostic and classification settings. In such problems, the underlying U-statistic is built from three independent multivariate samples, and the usual one- or two-sample JEL theory is not directly applicable. A general JEL framework for multivariate $3$-sample U-statistics is therefore needed.
The goal of this paper is to develop such a framework. We consider U-statistics built from three independent $q$-dimensional samples with degrees $(m_{1},m_{2},m_{3})$ and a symmetric kernel, and define a combined-sample construction that allows us to work with jackknife pseudo-values that are asymptotically independent but not identically distributed. Within this setting, we obtain a jackknife empirical likelihood ratio for a general scalar parameter $\theta$ defined through a three-sample U-statistic and establish a Wilks-type theorem: under mild moment and sample-size conditions, the JEL log-likelihood ratio converges in distribution to a chi-square random variable with one degree of freedom. This yields asymptotically valid confidence intervals for $\theta$ without explicit variance estimation or resampling.
To illustrate the practical relevance of the proposed methodology, we focus on problems arising in diagnostic accuracy assessment. For three-class classification, we construct confidence intervals for the difference between two correlated volumes under ROC surfaces (VUSs). In this case, the parameters of interest can be written as functionals of multivariate three-sample U-statistics, and our general JEL theory provides a coherent nonparametric inference procedure. Extensive Monte Carlo simulations under several multivariate models, including the Marshall–Olkin bivariate exponential and FGM-based bivariate Pareto distributions, are used to study the finite-sample performance of the proposed JEL intervals in comparison with intervals based on asymptotic normality with jackknife variance estimation and kernel-based bootstrap methods. Finally, we apply the methods to biomarker data from the Alzheimer’s Disease Neuroimaging Initiative, where we construct confidence intervals for differences in VUSs when comparing diagnostic markers across cognitive groups.
The paper is structured as follows. Section 2 we study $q$-dimensional U-statistics based on three samples. We study its asymptotic properties and find a consistent estimator of the unknown variance of the U-statistic based on Jackknife pseudo values. Jackknife empirical likelihood approach for these U-statistics is also discussed. As an illustration,in Section 3, we find confidence intervals for the difference between two dependent volumes under surfaces (VUS) using asymptotic normality of U-statistics, JEL method and bootstrapping of kernel type estimators . Section 4 presents extensive simulation studies assessing the performance of these confidence intervals under the Marshall–Olkin bivariate exponential and Bivariate Pareto models. Section 4 illustrates the methodology through a real-data analysis. Section 5 highlights the findings of the study along with some open problems.
2. Three sample multivariate U-Statistics
@@
Suppose $\left(\mathbb{X}_{1},\ldots,\mathbb{X}_{n_{1}}\right),\left(\mathbb{Y}_{1},\ldots,\mathbb{Y}_{n_{2}}\right)$ and $\left(\mathbb{Z}_{1},\ldots,\mathbb{Z}_{n_{3}}\right)$ represents three independent random samples from $q$-dimensional distribution functions $F$, $G$ and $H$, respectively, where $\mathbb{X}_{i}=(X_{i1},X_{i2},\dots,X_{iq})^{T},\;i=1,2,...,n_{1}$, $\mathbb{Y}_{j}=(Y_{j1},Y_{j2},\dots,Y_{jq})^{T},\;$ $j=1,2,...,n_{2},$ and $\mathbb{Z}_{k}=(Z_{k1},Z_{k2},\dots,Z_{kq})^{T},\;k=1,2,...,n_{3}$, and $q\geq 1$.
Let
$\theta=E[h\left(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}}\right)]\in\Re$
be the parameteric function of interest, where $h:\mathcal{X}^{m_{1}}\times\mathcal{Y}^{m_{2}}\times\mathcal{Z}^{m_{3}}\;\rightarrow\;\Re$ is a measurable function symmetric in its arguments, i.e., invariant under any permutation of indices within each sample, and $\mathcal{X},\mathcal{Y},$ and $\mathcal{Z}$ be the sample spaces associated with the distribution functions $F$, $G$, and $H$, respectively.
@@
An unbiased estimator of $\theta$ based on three-sample U-statistic with a symmetric kernel $h$ of degree $\left(m_{1},m_{2},m_{3}\right)$ is given by
@@
$U_{n_{1},n_{2},n_{3}}=$ $\frac{1}{{n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}\choose m_{3}}}\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ 1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{2}\\ 1\leq k_{1}<\ldots<k_{m_{3}}\leq n_{3}\end{subarray}}h\left(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}}\right),$ (2.1)
@@
2.1. Three sample multivariate U-Statistic -Asymptotic results
In this section we study the asymptotic properties of the U-statistic $U_{n_{1},n_{2},n_{3}}$. Define
@@ 2.1. Three sample multivariate U-Statistic -Asymptotic results
(a) $U_{n_{1},n_{2},n_{3}}\xrightarrow{\text{ a.s }}\theta$ as $\min\{n_{1},n_{2},n_{3}\}\rightarrow\infty$;
(b) Suppose that $\sigma_{100}^{2},\sigma_{010}^{2},\sigma_{001}^{2}>0$ holds. Let $\sigma^{2}=\frac{m_{1}^{2}\sigma_{100}^{2}}{n_{1}}+\frac{m_{2}^{2}\sigma_{010}^{2}}{n_{2}}+\frac{m_{3}^{2}\sigma_{001}^{2}}{n_{3}},$ Then,
$\frac{U_{n_{1},n_{2},n_{3}}-\theta}{\sigma}\xrightarrow{d}N(0,1),\text{ as }\min\{n_{1},n_{2},n_{3}\}\rightarrow\infty.$
For the proof of part (a) of the above theorem, see pp. 151–153 of korolyuk. The proof of part (b), i.e., the asymptotic normality of $U_{n_{1},n_{2},n_{3}}$, follows from puri1971nonparametric.
Since $\sigma^{2}$ is unknown, we find its consistent estimator. Following steps outline how to define a jackknife pseudo-value-based estimator. Consider
(1) the U statistics based on all observations $U_{n_{1},n_{2},n_{3}}$, as defined by (2.1);
$U_{n_{1}-1,n_{2},n_{3}}^{(-i)}=\frac{1}{{n_{1}-1\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}\choose m_{3}}}\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}-1\\ i_{l}\neq i,\;l=1,2,\dots,m_{1},\\ 1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{2}\\ 1\leq k_{1}<\ldots<k_{m_{3}}\leq n_{3}\end{subarray}}\thinspace h\left(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}}\right);$ (2.2)
$U_{n_{1},n_{2}-1,n_{3}}^{(-j)}=\frac{1}{{n_{1}\choose m_{1}}{n_{2}-1\choose m_{2}}{n_{3}\choose m_{3}}}\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ 1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{2}-1\\ j_{l}\neq j,\;l=1,2,\dots,m_{2},\\ 1\leq k_{1}<\ldots<k_{m_{3}}\leq n_{3}\end{subarray}}\thinspace h\left(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}}\right);$ (2.3)
$U_{n_{1},n_{2},n_{3}-1}^{(-k)}=\frac{1}{{n_{1}\choose m_{1}}{n_{2}\choose m_{2}}{n_{3}-1\choose m_{3}}}\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ 1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{2}\\ 1\leq k_{1}<\ldots<k_{m_{3}}\leq n_{3}-1\\ k_{l}\neq k,\;l=1,2,\dots,m_{3}\end{subarray}}\thinspace h\left(\mathbb{X}_{i_{1}},\ldots,\mathbb{X}_{i_{m_{1}}},\mathbb{Y}_{j_{1}},\ldots,\mathbb{Y}_{j_{m_{2}}},\mathbb{Z}_{k_{1}}\ldots,\mathbb{Z}_{k_{m_{3}}}\right).$ (2.4)
The jackknife pseudo-values for individual samples are defined as,
$V_{i,0,0}=n_{1}U_{n_{1},n_{2},n_{3}}-\left(n_{1}-1\right)U_{n_{1}-1,n_{2},n_{3}}^{(-i)},\;\;\forall\;\;i=1,2,\dots,n_{1},$ (2.5)
$V_{0,j,0}=n_{2}U_{n_{1},n_{2},n_{3}}-\left(n_{2}-1\right)U_{n_{1},n_{2}-1,n_{3}}^{(-j)},\;\;\forall\;\;j=1,2,\dots,n_{2}$ (2.6)
and
$V_{0,0,k}=n_{3}U_{n_{1},n_{2},n_{3}}-\left(n_{3}-1\right)U_{n_{1},n_{2},n_{3}-1}^{(-k)},\;\;\forall\;\;k=1,2,\dots,n_{3}.$ (2.7)
The jackknife pseudo-values based estimator of $\sigma^{2}$, proposed by arvesen1969jackknifing (also see wang2010empirical, guangming2013nonparametric and an2018jackknife),is given as
$+\frac{1}{n_{3}\left(n_{3}-1\right)}\sum_{k=1}^{n_{3}}\left(V_{0,0,k}-\bar{V}_{0,0,\cdot}\right)^{2},$ (2.8)
where $V_{i,0,0}$, $V_{0,j,0}$ and $V_{0,0,k}$ are defined in equations (2.5), (2.6) and (2.7), respectively.
Lemma 2.1.
Suppose that $\sigma_{100}^{2},\sigma_{010}^{2},\sigma_{001}^{2}>0$. Then
$\hat{\sigma}^{2}-\sigma^{2}=o_{p}\!\left(\min\{n_{1},n_{2},n_{3}\}^{-1}\right),$
where $\sigma^{2}=\frac{m_{1}^{2}\sigma_{100}^{2}}{n_{1}}+\frac{m_{2}^{2}\sigma_{010}^{2}}{n_{2}}+\frac{m_{3}^{2}\sigma_{001}^{2}}{n_{3}},$ and $\hat{\sigma}^{2}$ is defined in (2.1).
Proof: The proof can be easily derived using Chebyshev’s inequality and is therefore omitted (see guangming2013nonparametric).
Remark 2.2.
(i) From Theorem 2.1 we get that $U_{n_{1},n_{2},n_{3}}$ is a consistent estimator of $\theta$.
(ii) From Lemma 2.1 we get that $\hat{\sigma}^{2}$ is a consistent estimator of $\sigma^{2}$.
(iii) From Theorem 2.1 and Lemma 2.1,it follows that a two-sided $(1-\alpha)$ level asymptotic CI for $\theta$ is given by
@@
2.2. Three sample multivariate U-Statistics - JEL
jing2009jackknife showed that the JEL method, based on one and two-sample U-statistics, outperforms the normal approximation method. We study generalization of the JEL method to accommodate q-dimensional U-statistics arising from three samples. By broadening the applicability of JEL beyond conventional one and two-sample univariate settings, our approach enhances its relevance to vector valued three-sample statistical data analysis.
We combine the three samples and treat them as a single sample. For this purpose, let $n=n_{1}+n_{2}+n_{3}$ and define $\mathbb{W}_{l}$ as